Sterling Optical and Royal Optical both make glass frames and each is able to generate earnings before interest and taxes of $113,600. The separate capital structures for Sterling and Royal are shown here Sterling Royal Debt @ 8% Common stock, $5 $ 852.000 Debt @ 8% $ 284,000 1,136,000 $1,420,000 227,200 568,000 Common stock, $5 par Total Common shares $1,420,000 Total 113,600 Common shares a. Compute earnings per share for both firms. Assume a 30 percent tax rate. (Round your answers to 2 decimal places.) Earnings Share Sterling Royal b. In part a, you should have gotten the same answer for both companies earnings per share. Assuming a P/E ratio of 24 for each company, what would its stock price be? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Stock price c. Now as part of your analysis, assume the P/E ratio would be 18 for the riskier company in terms of heavy debt utilization in the capital structure and 22 for the less risky company. What would the stock prices for the two firms be under these assumptions? (Note: Although interest rates also would likely be different based on risk, we will hold them constant for ease of analysis.) (Do not round intermediate calculations. Round your answers to 2 decimal places.) Stock Price Sterling Royal
Answer #1
Particulars Sterling Royal
EBIT 113600 113600
LESS: Interest on debt 68160 19840
Earnings before TAX 45400 93760
Less tax @ 30% 13632 28128
Earnings after tax( EAT) 31808 65632

Number of shares of sterling is 113600

Therefore EPS = EAT/no.of shares

=31808/113600

=0.28

Number of shares of royal is 227200

EPS = 65632/ 227200

= 0.28

B.

To determine stlock price we use the formula EPS/PE ratio

=0.28/0.24

= 1.67

Is the stock price of both the companies.

C.

Sterling is the firm having more debt and is more riskier than the royal firm.

Therefore the PE ratio of sterling is 18% and for royal is 22%.

Stock price of sterling = EPS/ PE ratio

=0.28/0.18

= 1.55

Stock price of royal= 0.28/0.22

= 1.27